The supplement of an angle measures 40° more than twice its complement. Find the measure of the angle.
Answer by mananth(16946)

let the measure of angle be x
complement will be 90-x
40 more than twice its complement
= 2(90-x)+40
..
180-x is the supplement
180-x = 2(90-x)+40
180-x=180-2x+40
2x-x=40
x=40